OpenStudy (anonymous):

Calculate the integral. dz/z^2+36 from neg infinity to infinity. I got the the integral (1/6arctan(z/6)) and tried to solve it going from 0 to infinity but I just end up with 0. Am i missing something?

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty} dz/z^2+36\]

OpenStudy (anonymous):

nevermind, i figured it out. :>